OpenStudy (anonymous):

A snow cone stand sells crushed, flavored ice in cone-shaped paper cups. The opening of the cup is 8 cm wide and the height of the cup is 12cm . What is the approximate volume of the cup itself?

OpenStudy (anonymous):

\[Volume=\frac{ 1 }{ 3 }\pi r^2 h\] \[Volume = \frac{ 1 }{ 3 } \pi \times 4^2 \times 12 = 64\pi cm^3 \approx 201.1cm^3\]